# Using Dynamic Safety Margins as Control Barrier Functions

### v4

@@ 4 Lyapunov-based Dynamic Safety Margins
which implies that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq\Gamma(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, we must have $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma^{*}(\boldsymbol{\mathbf{v}})$. If $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, then $\boldsymbol{\mathbf{x}}=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ because $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$, recall that $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ with $\boldsymbol{\mathbf{x}}\in\mathcal{N}$. Furthermore, since $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}})$ by Assumption 4, we have that $\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}\neq 0$. So, there exists a sufficiently small scalar $\delta>0$ such that $\hat{\boldsymbol{\mathbf{x}}}=\boldsymbol{\mathbf{x}}-\delta\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a contradiction because
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$
@@
To prove property (22b), note that since $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$, and $\mathcal{X}\times\mathcal{V}$ is compact in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, we must only show that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$. Let $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\tilde{\mathcal{C}}$ be a sequence that converges to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are closed, it follows that $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Now we argue $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. For a contradiction, assume there is $i\in\{1,2\}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By continuity of $\Delta_{i}$ on $\mathbb{R}^{n}\times\mathbb{R}^{l}$, there exists an open neighborhood $\tilde{\mathcal{N}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ of $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ wherein $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By convergence of $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}$ to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, there exists $N\in\mathbb{N}$ such that $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{N}}$, which raises a contradiction because $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{C}}$ implies that $\Delta(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\geq 0$. It remains to show that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$. Begin by noting that for all $n\in\mathbb{N}$, $(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\in\tilde{\mathcal{D}}$ implies that $V(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\geq\gamma_{1}\big(\|\boldsymbol{\mathbf{x}}_{n}-\bar{x}(\boldsymbol{\mathbf{v}}_{n})\|\big)$. It follows by continuity of all functions involved in the inequality and by the order limit theorem that $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big(\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big)$. For a contradiction, let us assume that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\partial\tilde{\mathcal{D}}$, which implies that $\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and, therefore, $\overline{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. However, recall that we have shown $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. Thus, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq(1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})$. This yields
$\frac{1}{1-\epsilon}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\bar{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
@@
which is a contradiction because $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big(\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big)>0$. So, we must have that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ and we conclude that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, and therefore compact.
@@
To prove property (22c), let $i\in\{1,2\}$ be given and pick $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Then, by (33b),
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$ (38)
@@
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}$ can be used for the DSM-CBF policy $\beta$ given in (31). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
Corollary 4.29.
@@
If $\Delta$ is a Lyapunov-based DSM as in (37), then $\forall\alpha\in\mathscr{K}$, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$.
Proof 4.30.
@@
Let $\alpha\in\mathscr{K}$ and $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
$\dot{\Delta}_{i}\big(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\geq-\alpha\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big).$
Therefore, $\big(\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$.
Remark 4.31.
@@
When multiple constraints are considered, it is usually simpler to define multiple safety threshold values $\Gamma^{*}_{i}$ and stack them into the form of (37). As long as each one is differentiable, Theorem 4.27 can still be applied to obtain a DSM $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ with $n_{c}>2$. In the event either $\Gamma^{*}_{i}$ or $\overline{\Gamma}$ are not continuously differentiable, any continuously differentiable, lower-bounding function can be used in its place.
Remark 4.32.
@@
When choosing the initial virtual reference $\boldsymbol{\mathbf{v}}$ in the augmented system (24), it is worth noting that any $\boldsymbol{\mathbf{v}}_{0}\in\mathcal{V}$ such that $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\tilde{\mathcal{C}}$ is valid. A desirable choice is
$\boldsymbol{\mathbf{v}}_{0}=\underset{\boldsymbol{\mathbf{v}}\in\mathcal{V}}{\arg\min}~V(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}),$ (39)
@@
because it tends to maximize the value of $\Delta$. If the above problem is infeasible, then $\boldsymbol{\mathbf{x}}_{0}$ is not a safe starting condition for the designed DSM.
Remark 4.33.
@@
The proposed method effectively transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [32, 33, 34] and numerical [36] tools. It is also worth noting that one can always linearize the system around the current reference $v$ and use the Lyapunov function of the linearized system.
5 DSM-CBF Synthesis Approach Summary
@@
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor (ERG), are (vector-valued) CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [24] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (33) for the prestabilized system and evaluating the safety (34) and stability (36) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (31). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (31). Additional examples can be found in [42], where the approach presented in this paper is specialized to Euler–Lagrange systems.
6 Examples
@@ 6.1 Anthill system
While this function is discontinuous, it is bounded by the smooth function $\Gamma_{s}^{*}(v)=\max\left(\min\big(\Gamma^{*}(v),\frac{1}{4}\big),-\frac{1}{4}\right)$. Thus, by Theorem 4.27, $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{2}$ defined as
@@
is a DSM for $\pi$ and, by Theorem 3.9, a (vector-valued) CBF with respect to the augmented system.
@@
The simulation results for each approach are shown in Fig. 3. All approaches are able to enforce safety and reach the closest safe point to the desired reference but differ in performance and computational complexity. As expected, our DSM-CBF approach exhibits behavior that more closely resembles the nominal behavior than the ERG. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF approach with horizon $T=1$ exhibits almost equal performance to our DSM-CBF approach. However, it must be noted that its computational footprint is significantly higher than our DSM-CBF approach due to the need at each time step to predict the backup trajectory forward in time. This computational footprint can be reduced by considering shorter time horizons (e.g. $T=0.1$), but performance worsens as showin in Fig. 3. The computational overhead of the safety filter associated with each appraoch is showin in Table 1. Moreover, the backup CBF approach is less numerically robust, as evidenced by the response to a longer time horizon $T=10$. The reason the backup CBF fails for longer horizons is that it is numerically searching for the open boundary of the domain of attraction $x<v_{\max}$. Since a slight overshoot in the numerical integration leads to the divergence of the backup trajectory prediction, the backup CBF fails if it pushes too close to the boundary of the domain of attraction. Conversely, the DSM-based CBF avoids this issue by using the stability threshold $0.99\,\overline{\Gamma}(v)$, which yields a closed set with enough numerical leeway that small perturbations do not exit the domain of attraction of the prestabilizing controller. Our DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
Table 1: Safety Filter Compute Times
@@
Candidate CBF 4.81 5.02 - -
DSM-CBF (Proposed) 4.76 5.57 4.07 4.74
Backup CBF $T=0.1$ 5.95 6.02 9.04 10.12
Backup CBF $T=1.0$ 6.60 7.01 15.41 18.30
Backup CBF $T=5.0$ 15.03 20.16 24.51 32.62

### v5

@@ IV Lyapunov-based Dynamic Safety Margins
So, we must have $V(\bm{x},\bm{v})=\Gamma^{*}(\bm{v})$. If $V(\bm{x},\bm{v})=0$, then $\bm{x}=\bar{x}(\bm{v})\in\mathcal{X}_{\bm{v}}$ because $\bm{v}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\bm{v})=V(\bm{x},\bm{v})>0$, recall that $\mathcal{X}^{c}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}$ is open in $\mathbb{R}^{n}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}$ with $\bm{x}\in\mathcal{N}$. Furthermore, since $\bm{x}\in\mathcal{D}_{\bm{v}}\setminus\{\bar{x}(\bm{v})\}$ is a regular point of $V(\cdot,\bm{v})$ by Assumption 4, we have that $\frac{\partial V}{\partial\bm{x}}\neq 0$. So, there exists a sufficiently small scalar $\delta>0$ such that $\hat{\bm{x}}=\bm{x}-\delta\frac{\partial V}{\partial\bm{x}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\hat{\bm{x}},\bm{v})<V(\bm{x},\bm{v})$. So, $\Gamma^{*}(\bm{v})=V(\bm{x},\bm{v})$ is a contradiction because
$\Gamma^{*}(\bm{v})=\inf_{\bm{x}\in\mathcal{X}^{c}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})\leq V(\hat{\bm{x}},\bm{v})<V(\bm{x},\bm{v}).$
@@
To prove property (19b), note that since $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$, and $\mathcal{X}\times\mathcal{V}$ is compact in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, we must only show that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$. Let $\{(\bm{x}_{k},\bm{v}_{k})\}\subset\tilde{\mathcal{C}}$ be a sequence that converges to $(\bm{x},\bm{v})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are closed, it follows that $\bm{x}\in\mathcal{X}$ and $\bm{v}\in\mathcal{V}$. Now we argue $\Delta(\bm{x},\bm{v})\geq 0$. For a contradiction, assume there is $i\in\{1,2\}$ such that $\Delta_{i}(\bm{x},\bm{v})<0$. By continuity of $\Delta_{i}$ on $\mathbb{R}^{n}\times\mathbb{R}^{l}$, there exists an open neighborhood $\tilde{\mathcal{N}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ of $(\bm{x},\bm{v})$ wherein $\Delta_{i}(\bm{x},\bm{v})<0$. By convergence of $\{(\bm{x}_{k},\bm{v}_{k})\}$ to $(\bm{x},\bm{v})$, there exists $K\in\mathbb{N}$ such that $(\bm{x}_{K},\bm{v}_{K})\in\tilde{\mathcal{N}}$, which raises a contradiction because $(\bm{x}_{K},\bm{v}_{K})\in\tilde{\mathcal{C}}$ implies that $\Delta(\bm{x}_{K},\bm{v}_{K})\geq 0$. It remains to show that $(\bm{x},\bm{v})\in\tilde{\mathcal{D}}$. Begin by noting that for all $k\in\mathbb{N}$, $(\bm{x}_{k},\bm{v}_{k})\in\tilde{\mathcal{D}}$ implies that $V(\bm{x}_{k},\bm{v}_{k})\geq\gamma_{1}\big(\|\bm{x}_{k}-\bar{x}(\bm{v}_{k})\|\big)$. It follows by continuity of all functions involved in the inequality and by the order limit theorem that $V(\bm{x},\bm{v})\geq\gamma_{1}\big(\|\bm{x}-\bar{x}(\bm{v})\|\big)$. For a contradiction, let us assume that $(\bm{x},\bm{v})\in\partial\tilde{\mathcal{D}}$, which implies that $\bm{x}\in\partial\mathcal{D}_{\bm{v}}$ and, therefore, $\overline{\Gamma}(\bm{v})\leq V(\bm{x},\bm{v})$. However, recall that we have shown $\Delta(\bm{x},\bm{v})\geq 0$. Thus, $V(\bm{x},\bm{v})\leq(1-\epsilon)\overline{\Gamma}(\bm{v})$. This yields
$\frac{1}{1-\epsilon}V(\bm{x},\bm{v})\leq\bar{\Gamma}(\bm{v})\leq V(\bm{x},\bm{v}),$
@@
which is a contradiction because $V(\bm{x},\bm{v})\geq\gamma_{1}\big(\|\bm{x}-\bar{x}(\bm{v})\|\big)>0$. So, we must have that $(\bm{x},\bm{v})\in\tilde{\mathcal{D}}$ and we conclude that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, and therefore compact.
@@
To prove property (19c), let $i\in\{1,2\}$ be given and pick $(\bm{x},\bm{v})\in\tilde{\mathcal{D}}$ such that $\Delta_{i}(\bm{x},\bm{v})=0$. Then, by (30b),
$\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})=-\frac{\partial V}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})\geq 0.$ (35)
@@
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}$ can be used for the DSM-CBF policy $\tilde{\beta}$ given in (28). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
Corollary 3
@@
If $\Delta$ is a Lyapunov-based DSM as in (34), then $\forall\alpha\in\mathscr{K}$, $\forall(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})\neq\emptyset$.
Proof:
@@
Let $\alpha\in\mathscr{K}$ and $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\bm{x},\bm{v})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
$\dot{\Delta}_{i}\big(\bm{x},\bm{v},\pi(\bm{x},\bm{v}),0\big)=-\frac{\partial V}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})\geq 0\geq-\alpha\big(\Delta_{i}(\bm{x},\bm{v})\big).$
Therefore, $\big(\pi(\bm{x},\bm{v}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})$. ∎
Remark 9
@@
When multiple constraints are considered, it is usually simpler to define multiple safety threshold values $\Gamma^{*}_{i}$ and stack them into the form of (34). As long as each one is differentiable, Theorem 7 can still be applied to obtain a DSM $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ with $n_{c}>2$. In the event either $\Gamma^{*}_{i}$ or $\overline{\Gamma}$ are not continuously differentiable, any continuously differentiable, lower-bounding function can be used in its place.
Remark 10
@@
When choosing the initial virtual reference $\bm{v}$ in the augmented system (21), it is worth noting that any $\bm{v}_{0}\in\mathcal{V}$ such that $(\bm{x}_{0},\bm{v}_{0})\in\tilde{\mathcal{C}}$ is valid. A desirable choice is
$\bm{v}_{0}=\underset{\bm{v}\in\mathcal{V}}{\arg\min}~V(\bm{x}_{0},\bm{v}),$ (36)
@@
because it tends to maximize the value of $\Delta$. If the above problem is infeasible, then $\bm{x}_{0}$ is not a safe starting condition for the designed DSM.
Remark 11
@@
This section transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [37, 24, 39] and numerical [16] tools. It is also worth noting that one can always linearize the system around the current reference $v$ and use the Lyapunov function of the linearized system.
V DSM-CBF Synthesis Approach Summary
@@
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor (ERG), are CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [38] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (30) for the prestabilized system and evaluating the safety (31) and stability (33) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (28). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (28). Additional examples can be found in [20], where the approach presented in this paper is specialized to Euler–Lagrange systems.
VI Examples
@@ VI-A Anthill system
While this function is discontinuous, it is bounded by the smooth function $\Gamma_{s}^{*}(v)=\max\left(\min\big(\Gamma^{*}(v),\frac{1}{4}\big),-\frac{1}{4}\right)$. Thus, by Theorem 7, $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{2}$ defined as
@@
is a DSM and, by Theorem 4, a CBF with respect to the augmented system.
@@
The simulation results for each approach are shown in Fig. 3. All approaches are able to enforce safety and reach the closest safe point to the desired reference but differ in performance and computational complexity. As expected, our DSM-CBF approach exhibits behavior that more closely resembles the nominal behavior than the ERG. With enough tuning, the candidate CBFs were able to also maintain safety for the fist 10 seconds, albeit with poor performance and no guarantee they will remain feasible afterwards. This is unsurprising since the candidate CBFs were constructed without any understanding of the inherent control authority constraint. The backup CBF approach with horizon $T=1$ exhibits almost equal performance to our DSM-CBF approach. However, it must be noted that its computational footprint is significantly higher than our DSM-CBF approach due to the need at each time step to predict the backup trajectory forward in time. This computational footprint can be reduced by considering shorter time horizons (e.g. $T=0.1$), but performance worsens as showin in Fig. 3. The computational overhead of the safety filter associated with each appraoch is showin in Table I. Moreover, the backup CBF approach is less numerically robust, as evidenced by the response to a longer time horizon $T=10$. The reason the backup CBF fails for longer horizons is that it is numerically searching for the open boundary of the domain of attraction $x<v_{\max}$. Since a slight overshoot in the numerical integration leads to the divergence of the backup trajectory prediction, the backup CBF fails if it pushes too close to the boundary of the domain of attraction. Conversely, the DSM-based CBF avoids this issue by using the stability threshold $0.99\,\overline{\Gamma}(v)$, which yields a closed set with enough numerical leeway that small perturbations do not exit the domain of attraction of the prestabilizing controller. Our DSM-based CBF provided a systematic approach that successfully identified the constraints induced by the input saturation.
TABLE I: Safety Filter Compute Times
@@
Candidate CBF 3.46 4.27 - -
DSM-CBF (Proposed) 3.54 4.76 3.64 6.94
Backup CBF $T=0.1$ 4.51 5.33 7.85 15.26
Backup CBF $T=1.0$ 4.99 6.11 12.98 22.93
Backup CBF $T=5.0$ 10.92 17.25 20.39 35.97